OpenStudy (baby456):

Define a piecewise function basedon the description below please help me write a piecewise function based on the description below. A) You determine your tax credit,C, based an annual salary,a. If your annual salary is below 24,000$, your tax credit is based on 25% of the salry. if the salary is between 24,000 and 50,000 or more, the credits is 14% of the annual salary.

OpenStudy (baby456):

@satellite73 @jim_thompson5910 @zepdrix @zzr0ck3r

OpenStudy (danjs):

ok, so this has 2 different conditions, if you make 0 to 24k and if you make 24k and up, the tax amount is different..

OpenStudy (danjs):

Say you make 'a' dollars, the tax credit is C(a) for 25%, C(a) = (0.25)*a for 14%, C(a) = (0.14)*a

OpenStudy (danjs):

 one for 0 to 24000, and one for greater than 24000